the length of a rectangle is 3 yd less than twice its width, and the area of the rectangle is 27^2. find the dimensions
Answer by Cromlix(4381)

Width = x
Length = 2x - 3
x(2x - 3) = 27
2x^2 - 3x - 27 = 0
(2x - 9)(x + 3) = 0
2x - 9 = 0
2x = 9
x = 4.5 yds
x + 3 = x = -3(no answer)
Width = 4.5 yards
Length = 6 yards (2(4.5) - 3)
Hope this helps.
:-)